### v2

# Test-negative designs with various reasons for testing: statistical bias and solution

Abstract
Test-negative designs are widely used for post-market evaluation of vaccine effectiveness, particularly in cases where randomization is not feasible. Differing from classical test-negative designs where only healthcare-seekers with symptoms are included, recent test-negative designs have involved individuals with various reasons for testing, especially in an outbreak setting. While including these data can increase sample size and hence improve precision, concerns have been raised about whether they introduce bias into the current framework of test-negative designs, thereby demanding a formal statistical examination of this modified design. In this article, using statistical derivations, causal graphs, and numerical simulations, we show that the standard odds ratio estimator may be biased if various reasons for testing are not accounted for. To eliminate this bias, we identify three categories of reasons for testing, including symptoms, disease-unrelated reasons, and case contact tracing, and characterize associated statistical properties and estimands. Based on our characterization, we show how to consistently estimate each estimand via stratification. Furthermore, we describe when these estimands correspond to the same vaccine effectiveness parameter, and, when appropriate, propose a stratified estimator that can incorporate multiple reasons for testing and improve precision. The performance of our proposed method is demonstrated through simulation studies.
Keywords: Covid-19, stratification, precision, vaccine effectiveness.
Introduction
The test-negative design (TND, [1]) refers to a study where patients presenting at healthcare facilities are tested for a disease of interest and categorized as test-positive or test-negative. The intervention effect of interest is then estimated as the odds ratio of intervention comparing test-positive and test-negative patients. As an observational study design, the TND has been increasingly used to study vaccine effectiveness (VE) for infectious diseases such as influenza [2], coronavirus [3], and rotavirus [4], especially when randomized trials are infeasible or unethical. Compared to a cohort design, the TND has the advantage of efficiency and cost-effectiveness, and can potentially correct for bias arising from differential healthcare-seeking behaviors [5].
Classical TNDs only recruit patients who actively seek healthcare and get tested for infection because of disease-related symptoms, but, in many recent applications, recruitment also enrolls patients with other reasons for testing, particularly apparent in Covid-19 vaccine studies. For example, in a TND study evaluating the effectiveness of face masks for preventing Covid-19 [6], only 36.4% patients were tested due to symptoms, accounting for 77.9% of the positive tests but only 16.7% of the negative tests. Other reasons for testing included routine screening through work or school (29.9%), testing required for medical procedures (12.4%), out of curiosity (12.2%), and travel screening (8%). In two other TNDs of Covid-19 that reported reasons for testing [7, 8], symptoms took a slightly smaller proportion of tested subjects (25.3% and 29.5%), while contact with positive cases emerged as a major reason for testing (52% and 18%). In addition to these examples, many Covid-19 TND studies involved asymptomatic patients [9], mandatory testing for healthcare workers [10], and massive community screening [11], none of which can be interpreted as testing due to symptoms. Beyond Covid-19, TNDs were also suggested for evaluating VE in an Ebola outbreak [12, 13], where some participants can be identified and tested due to close contact with known cases. Overall, such changes in patient inclusion criteria in TNDs often occur in an outbreak setting, where a great number of tests are not from symptom-related spontaneous healthcare seeking, and these tests are often too frequent to be overlooked.
Despite the emergence of TNDs that include various reasons for testing, appropriate methods to analyze such studies remain an open question. The status quo is to ignore the reasons for testing and perform a standard analysis, e.g., via odds ratios and logistic regression, aggregating the data across all reasons for testing. However, this method can be biased for estimating the overall VE since different reasons for testing may involve heterogeneous populations, across which the VE, the likelihood of testing, and covariate distributions can all differ. [14] pointed out that reasons other than symptoms for testing may introduce bias, and [15] conceptualized this issue through a causal graph. Both papers and [16] stated that a stratified analysis based on testing reasons could alleviate the bias, but no complete quantification or formal solution was given. [13] addressed the bias from a naive TND analysis when there were two reasons for testing: (i) symptomatic self-reporting or (ii) testing close contacts of known cases, and proposed a hybrid design to remove such bias. However, they made a strong assumption that the VE for either recruitment population is the same; in addition, other testing reasons, such as routine screening, were not considered in their setting. Finally, a naive solution is to filter out asymptomatic subjects and only use the remaining data to perform a classical TND analysis. Although this procedure is valid, it loses a large amount of information.
In this paper, we combine directed acyclic graphs (DAGs, [17]) with additional assumptions to examine the statistical properties of TNDs with various reasons for testing, and propose several estimands with corresponding unbiased estimators. First, we revisit the classical TND to clarify underlying assumptions and an appropriate estimand, and then identify two additional categories of reasons for testing: disease-unrelated reasons and contact tracing, for which we construct corresponding DAGs and derive the target estimands. The target estimands for these three categories are generally distinct, which expand the classical scope of TNDs and provide new perspectives on the generalizability of TND estimates. Second, given multiple reasons for testing, we explain why the odds ratio estimator pooling all testing reasons may be biased. Finally, we propose several estimators that can target each VE of interest using the corresponding data. Furthermore, we provide conditions to justify a combined estimator based on multiple reasons for testing, thereby maximizing the use of relevant data to achieve unbiased estimation, valid hypothesis testing, and increased precision. Moreover, our proposed methods do not solely depend on parametric model assumptions; they also offer the flexibility to incorporate non-parametric kernel smoothing to approximate the underlying data-generating distribution. To the best of our knowledge, we provide the first vehicle to clarify to what extent non-symptomatic reasons for testing can be included in TNDs without introducing bias, and our proposed methods are the first to accommodate a range of reasons for testing, which can inform statistical practice for future TNDs.
Defintions and assumptions
We consider TNDs to study a specific disease, namely Covid-19, with a view to estimating vaccine effectiveness. For simplicity, we consider assessment of only a single vaccination dose, although the development can be extended to cover situations where $V$ captures a more general vaccination history. Thus, for each subject, let $V$ be the vaccination status with $V=1$ if vaccinated and 00 otherwise. Let $I$ be the true infection status for severe acute respiratory syndrome coronavirus 2 (SARS-CoV-2), where $I=1$ if the subject is infected and $I=0$ otherwise, as determined by a definitive diagnostic test. Let $R$ be a categorical variable encoding the reasons for testing with mutually exclusive categories, with each category associated with a single reason, such as “routine screening for work”, or a pool of related reasons, such as “Covid-19-like symptoms”. For a traditional test-negative design, we let $T$ be a binary indicator of being tested at a healthcare setting (and thus eligible for inclusion in a standard test-negative study) with $T=1$ if tested and 00 otherwise. Throughout, we assume all the above variables are measured without misclassification. Finally, $\boldsymbol{X}$ denotes observed covariates, such as age, sex, and calendar time, and $H$ is a latent (unobserved) healthcare-seeking behavior indicator. Again, for simplicity, we assume $H$ is a binary random variable with $H=1$ indicating healthcare-seeking behavior for a symptomatic participant and $H=0$ otherwise. This setup also applies to studies on other infectious diseases, such as influenza, and other non-vaccine interventions with the only difference being that the reasons for testing may be different.
In a standard TND, all eligible patients with Covid-19-like symptoms who are tested are included by design and subsequently classified by whether they test positive for the disease in question—here infection with SARS-CoV-2—or not. This test is assumed to have perfect sensitivity and specificity. This enrollment scheme results in an observational study [1, 18]; the sampling scheme can be considered as a historical cohort study (as vaccination has occurred in the past) or as a form of case-cohort sampling.
A primary goal is inference on the conditional VE for strata encoded by observed covariates, which is simply one minus the conditional relative risk of infection given observed covariates,
$VE(\boldsymbol{x})=1-\dfrac{\Pr(I=1|V=1,\boldsymbol{X}=\boldsymbol{x})}{\Pr(I=1|V=0,\boldsymbol{X}=\boldsymbol{x})}.$ (1)
Of note, the conditional VE is a function of $\boldsymbol{x}$, which means inference is made on each stratum, such as a specific age group. VE across different strata can be linked by assuming, for example, that $\Pr(I=1|V,\boldsymbol{X})$ follows a log-binomial regression model
$\log\{\Pr(I=1|V,\boldsymbol{X})\}=\beta_{0}+\beta_{V}V+\beta_{\boldsymbol{X}}^{\top}\boldsymbol{X},$
where $\beta_{0},\beta_{V},\beta_{\boldsymbol{X}}$ are regression coefficients. Under this model, $VE(\boldsymbol{x})=1-e^{\beta_{V}}$ does not vary by $\boldsymbol{x}$, which can be interpreted as reflecting homogeneous VE across strata.
At first glance, $H$ is of less obvious interest for conditional VE estimation. However, $H$ may be associated with both $V$ and $I$, and in fact, TND studies were originally conceived as a way of eliminating, or at least reducing, this potential confounding bias by focusing only on the stratum with $H=1$. Thus, another estimand of interest is the conditional VE for strata defined by observed covariates $\boldsymbol{X}$ and the latent healthcare-seeking behavior $H$, defined as
$VE(\boldsymbol{x},h)=1-\dfrac{\Pr(I=1|V=1,\boldsymbol{X}=\boldsymbol{x},H=h)}{\Pr(I=1|V=0,\boldsymbol{X}=\boldsymbol{x},H=h)}.$ (2)
Compared to $VE(\boldsymbol{x})$ defined in Equation (1), $VE(\boldsymbol{x},h)$ further depends on the latent variable $H$ and is hence more challenging to estimate. As noted, the standard TND study automatically stratifies on $H$ as no data is observed for those with $H=0$. The two estimands $VE(\boldsymbol{x})$ and $VE(\boldsymbol{x},h)$ are connected through the collapsibility of relative risks [19]. Mathematically, if we assume healthcare-seeking behavior is not an effect modifier, i.e., $VE(\boldsymbol{x},1)=VE(\boldsymbol{x},0)$, and $V$ is independent of $H$ given $\boldsymbol{X}$, then $VE(\boldsymbol{x})=VE(\boldsymbol{x},h)$.
The two VE estimands characterize associations instead of causality. We next consider their links to causal estimands under structural causal assumptions.
Connection to casual interpretation
We define $I(v)$ as the potential outcome of a subject if their vaccination status were $V=v$ for $v=0,1$. For a causal interpretation of $VE(\boldsymbol{x},h)$, we define
[…]
To provide a causal interpretation for $VE(\boldsymbol{x})$, we have two approaches. The first approach assumes that healthcare-seeking is not a confounder to obtain
${VE(\boldsymbol{x})=\textup{causal-}VE(\boldsymbol{x})=1-\dfrac{\Pr(I(1)=1\mid\boldsymbol{X}=\boldsymbol{x})}{\Pr(I(0)=1\mid\boldsymbol{X}=\boldsymbol{x})}}$
under the above causal assumptions for $VE(\boldsymbol{x},h)$. A second approach assumes the same causal assumptions for $VE(\boldsymbol{x},h)$ and uses proximal learning [21] to address the issue of unmeasured confounders by leveraging measured proxy variables such as vaccine history.
A concise review of statistical analyses for classical TNDs
To illustrate the underlying causal structure of TNDs, we use directed acyclic graphs (DAGs, [17]), which describe the relationships (arrows) among variables (nodes) and are widely used in medical and epidemiological research. Of note, the DAG itself does not encode all necessary assumptions for TNDs [22], and we combine DAGs and additional assumptions for our derivation (See Table 1). Figure 1 shows a DAG for classical TNDs where infection-related symptoms are the only reason for testing. Here, we additionally denote $S$ as an indicator of having Covid-19-like symptoms, with $I=1$ and $S=1$ together implying being diseased. By definition, $R=\textup{Symptoms}$ implies $S=1$. We assume that patients cannot self-diagnose based on symptoms, which is reflected in Figure 1 by no $I\rightarrow R=\textup{Symptoms}$ or $I\rightarrow T$ direct effects. In addition, an important feature of classical TNDs is that enrolled individuals are all healthcare-seekers who complete the test and have Covid-19-like symptoms, i.e., those with $H=1$, $S=1$, and $T=1$, thereby providing no information on VE among healthcare non-seekers or untested individuals.
$\boldsymbol{X}$$V$$I$$S=1$ $R=$ Symptoms $T=1$$H=1$
Figure 1: The DAG for classical TNDs. Circle nodes stand for random variables, while rectangle nodes represent that the random variables are conditioned on specific values.
Beyond the DAG, we assume that vaccination does not reduce non-SARS-CoV-2 causes of symptoms among healthcare-seekers [1], i.e.,
${\Pr(I=0,S=1|V=1,\boldsymbol{X},H=1)=\Pr(I=0,S=1|V=0,\boldsymbol{X},H=1).}$
These assumptions together motivate an odds-ratio estimator for the VE on the relative risk scale:
[…]
$={1-\frac{\Pr(I=1,S=1|V=1,\boldsymbol{X},H=1)}{\Pr(I=1,S=1|V=0,\boldsymbol{X},H=1)}.}$ (3)
We denote the right-hand side of Equation (3) as $VE_{s}(\boldsymbol{X},1)$, as it evaluates VE for symptomatic infections among healthcare-seekers. Here, standard TNDs achieve VE identification via odds ratio because of its unique design that controls are also healthcare seekers with Covid-like symptoms and that vaccination has no effect on other types of infections, which together cancels out the collider bias from $T$.
In addition, $VE_{s}(\boldsymbol{X},1)$ may differ from $VE(\boldsymbol{X},1)$ since $VE(\boldsymbol{X},1)$ also accounts for vaccination effects on the risk of asymptomatic infections. This relationship can be made explicit mathematically. Specifically, if $RR_{P}$ is the relative risk of progression to the defined symptom criteria conditional on infection, comparing vaccinated to unvaccinated individuals, then
${VE_{s}(\boldsymbol{X},1)=1-[1-VE(\boldsymbol{X},1)]RR_{P}}.$
The difference between $VE(\boldsymbol{X},1)$ and $VE_{s}(\boldsymbol{X},1)$ thus depends on $RR_{P}$, which in turn is potentially influenced by the definition of symptoms required for recruitment.
To require equivalence of $VE_{s}(\boldsymbol{X},1)$ and $VE(\boldsymbol{X},1)$, we thus need an additional restrictive assumption that vaccination has no effect on preventing symptom development given infection, i.e., $RR_{P}=1$ or, equivalently,
${\Pr(S=1|V=1,I=1,\boldsymbol{X},H=1)=\Pr(S=1|V=0,I=1,\boldsymbol{X},H=1).}$
Although this assumption is known to be violated for Covid-19 vaccines, it may be plausible if $V$ represents non-vaccine interventions, e.g., wearing facial masks [6], social distancing, or other hygiene behaviors. Given this assumption, classical TNDs can target $VE(\boldsymbol{x},1)$. However, vaccination often leads to symptom alleviation so that $VE_{s}(\boldsymbol{x},1)$ is usually larger than $VE(\boldsymbol{x},1)$, indicating increased effectiveness for preventing symptomatic infections.
Two additional categories of reasons for testing
We next consider two common types of reasons for testing other than symptoms and discuss their target VE. We name the two categories of reasons as “disease-unrelated reasons” and “case contact tracing”, each of which is described with a DAG as presented in Figures 2(a) and 2(b), respectively. For these two categories, we assume a cohort sampling scheme, which better reflects the actual recruitment procedure than case-cohort sampling and facilitates the statistical identification.
$\boldsymbol{X}$$H$$V$$I$$S$ $R=$ Disease- unrelated reasons $T=1$
$\boldsymbol{X}$$H$$V$$I$$S$ $R=$ Case contact tracing $T=1$
Figure 2: DAGs for testing due to disease-unrelated reasons (a) and case contact tracing (b). Circle nodes stand for random variables, while rectangle nodes represent that the random variables are conditioned on specific values.
[…]
$VE(\boldsymbol{x})=1-\frac{\Pr(I=1|V=1,\boldsymbol{X}=\boldsymbol{x},R=\textrm{Disease-unrelated reasons},T=1)}{\Pr(I=1|V=0,\boldsymbol{X}=\boldsymbol{x},R=\textrm{Disease-unrelated reasons},T=1)}$
As a result, the relative risk obtained from tested individuals with disease-unrelated reasons targets $VE(\boldsymbol{x})$, covers both healthcare-seekers and non-seekers, and directly assesses vaccine effects on infections without any assumption but effects on symptoms.
Case contact tracing is a common practice of monitoring the spread of diseases in an outbreak setting by testing all contacts of known confirmed cases. Because of the mandatory nature of such contact tracing, testing is unrelated to healthcare-seeking behavior or vaccination status, but it can be related to covariates (e.g., occupation), and contacts of known cases may be more likely to suffer infection and symptoms ($R\rightarrow I$ and $R\rightarrow S$ direct effects in Figure 2(b)). Given these assumptions, we define
$VE_{cct}(\boldsymbol{x})$ $=1-\frac{\Pr(I=1|V=1,\boldsymbol{X}=\boldsymbol{x},R=\textrm{Casecontact tracing},T=1)}{\Pr(I=1|V=0,\boldsymbol{X}=\boldsymbol{x},R=\textrm{Case contact tracing},T=1)}$
$=1-\frac{\Pr(I=1|V=1,\boldsymbol{X}=\boldsymbol{x},R=\textrm{Casecontact tracing})}{\Pr(I=1|V=0,\boldsymbol{X}=\boldsymbol{x},R=\textrm{Case contact tracing})},$ (4)
representing the VE among contacts of known cases.
To connect $VE_{cct}(\boldsymbol{x})$ with the target $VE(\boldsymbol{x})$, we need an additional but reasonable assumption that being a contact of known cases does not modify VE, i.e.,
$\frac{\Pr(I=1|V=1,\boldsymbol{X},R=\textrm{Case contact tracing})}{\Pr(I=1|V=0,\boldsymbol{X},R=\textrm{Case contact tracing})}=\frac{\Pr(I=1|V=1,\boldsymbol{X},R\neq\textrm{Case contact tracing})}{\Pr(I=1|V=0,\boldsymbol{X},R\neq\textrm{Case contact tracing})}.$
This assumption implies $VE_{cct}(\boldsymbol{x})=VE(\boldsymbol{x})$ so that we are able to use the tested case contact subjects to estimate $VE(\boldsymbol{x})$. In practice, close contacts of cases might represent a higher level of exposure intensity to the virus, a potential confounder for VE. Therefore, case contact tracing may cause VE modification if the exposure intensity is not measured, in which cases $VE_{cct}(\boldsymbol{x})\neq VE(\boldsymbol{x})$. On the other hand, whether case contact tracing indicates a higher level of exposure intensity (compared to infection in the general population) needs further justification. This is because some contacts in contact tracing may be casual contacts, and general infections may result from close contacts, perhaps in their household.
Overall, Table 1 summarizes the reasons for testing, their assumptions, and the target VE discussed above. Under minimum assumptions, each category of reasons for testing targets its own VE estimand. In addition, classical TNDs are constrained to healthcare seekers, whereas “disease-unrelated reasons” and “case contact tracing” cover both healthcare-seekers and non-seekers. Categorization of reasons for testing thus demonstrates the bias from an aggregate odds ratio estimator and also provides a stepping stone for strategically combining reasons for testing, which we introduce below.
Table 1: A summary of considered reasons for testing with their assumptions, examples, and target VEs in TNDs. $VE(\boldsymbol{x})$, $VE(\boldsymbol{x},1)$, $VE_{s}(\boldsymbol{x},1)$ and $VE_{cct}(\boldsymbol{x})$ are defined in Equations (1), (2), (3), (4), representing the conditional VE for infection, conditional VE for infection among healthcare-seekers, conditional VE for symptomatic infection (and testing due to symptoms), and conditional VE for infection among contacts of known cases.
Reasons for testing Main assumptions Examples Target VEs
Symptoms (Figure 1) 1. Only healthcare-seekers are tested. 2. Vaccination does not reduce     test-negative symptoms. 3. Given a symptom, the uptake of testing     is the same among cases (test-positives)     and controls (test-negatives). 4. Vaccination has no direct effect     on symptom severity among the infected. Fever, fatigue $VE(\boldsymbol{x},1)$; or $VE_{s}(\boldsymbol{x},1)$ without assumption 4.
Disease- unrelated reasons (Figure 2(a)) 1. Testing is not directly impacted     by infection status or     healthcare-seeking behavior. Surveillance tests, travelling $VE(\boldsymbol{x})$
Case contact tracing (Figure 2(b)) 1. Testing is not directly impacted by     healthcare-seeking behavior. 2. Contact of known cases is not an effect     modifier. Case contact tracing $VE(\boldsymbol{x})$; or $VE_{cct}(\boldsymbol{x})$ without assumption 2.
The above framework also works if the outcome of interest is symptomatic infection, i.e., $I=1$ and $S=1$. With this new outcome definition, testing due to symptoms as in classical TNDs naturally targets $VE_{s}(\boldsymbol{x},1)$ as described in Equation (3). For tests due to disease-unrelated reasons and case contact tracing, the relative risks can be updated by focusing on $I=1,S=1$; this change will require knowing whether tested individuals have symptoms, which may be less realistic in many scenarios. For example, certain prescribed medications may prevent or delay symptom development, testing may block symptoms occurring (particularly if symptoms are severe, e.g., require hospitalization), and symptom evaluation may be unavailable during testing.
Potential bias given various reasons for testing
$\boldsymbol{X}$$H$$V$$I$$S$$R$$T=1$
Figure 3: The DAG for TNDs with multiple testing reasons. Circle nodes stand for random variables, while rectangle nodes represent that the random variables are conditioned on specific values.
When a TND involves multiple reasons for testing (including the above-discussed three categories or other reasons), different causal relationships occur, as demonstrated in Figure 3. For example, when some individuals are tested for onsite work, test results arise from both healthcare-seekers and non-seekers. When testing sometimes occurs due to Covid-specific symptoms (e.g. loss of sense of smell), this may lead to a direct relationship $I\rightarrow T$. Also, in some cases, asymptomatic testing is more likely to occur in healthcare-seekers, leading to a direction relationship $H\rightarrow R$. If such multiple testing reasons all occur in the data, we will obtain a DAG akin to the one shown in Figure 3, where all variables are fully connected. In this case, a naive odds ratio estimator with observed data will be biased for VE, i.e.,
[…]
since the collider bias arising from conditioning on $T=1$ cannot be removed. This bias also exists for estimating $VE(\boldsymbol{x},h)$ for similar reasons and, additionally, due to the unmeasured confounder $H$. As a numerical example, consider a binary confounder $X$ with $\Pr(X=1)=0.5$, $\Pr(V=1|X,H)=0.5$, and $\log\{\Pr(I=1|V,X,H)\}=-0.5-0.5V-1.5X$. We consider testing due to symptoms, screening, or other reasons with
$\Pr(T=1,R=r|I,V,X,H)=\left\{\begin{array}[]{rl}0.2X+0.2IX&\textrm{if}\ r=\textrm{Symptoms},\\ 0.4V&\textrm{if}\ r=\textrm{Screening},\\ 0.1VI&\textrm{if}\ r=\textrm{Other}.\end{array}\right.$
Then, the odds ratio VE estimator based on aggregated results from tested individuals is 0.02, which substantially underestimates the VE among the whole population, which is 0.39. In addition, we have $VE(\boldsymbol{x})=VE(\boldsymbol{x},1)=VE_{s}(\boldsymbol{x},1)=0.39$ in this example, which demonstrates that the aggregated odds ratio estimator may not target any VE of interest.
Unbiased estimation by stratification
For a TND with multiple reasons for testing, we next show how to estimate the identifiable VEs (summarized in Table 1), based on each type of reason
[…]
$\Pr(I=1|V,\boldsymbol{X},R=r,T=1)=h(V,\boldsymbol{X},r),$
where $h(V,\boldsymbol{X},r)$ is a user-specified learning model. We can assume $h(V,\boldsymbol{X},r)$ follows a parametric model, such as a log-binomial regression model, or a non-parametric model, using kernel smoothing for non-linear effects of $X$; both approaches are well-developed in the literature with software available for estimation. Compared to parametric models, a non-parametric model requires fewer assumptions and can better approximate the truth with the goal of reducing bias, albeit with the disadvantages of requiring larger sample sizes to avoid overfitting and increased variance associated with a slower convergence rate. We provide mathematical details and model fitting in Web Appendix B. With a user-specified model, we obtain $\widehat{h}(V,\boldsymbol{X},r)$ as the estimate of $h(V,\boldsymbol{X},r)$.
We next construct the following estimators using data from each pre-defined category of reasons for testing:
$\widehat{VE}_{s}(\boldsymbol{x},1)$ $=1-\frac{\widehat{h}(V=1,\boldsymbol{X}=\boldsymbol{x},r=\textup{Symptoms})\{1-\widehat{h}(V=0,\boldsymbol{X}=\boldsymbol{x},r=\textup{Symptoms})\}}{\widehat{h}(V=0,\boldsymbol{X}=\boldsymbol{x},r=\textup{Symptoms})\{1-\widehat{h}(V=1,\boldsymbol{X}=\boldsymbol{x},r=\textup{Symptoms})\}},$
$\widehat{VE}(\boldsymbol{x})$ $=1-\frac{\widehat{h}(V=1,\boldsymbol{X}=\boldsymbol{x},r=\textup{Disease-unrelated reasons})}{\widehat{h}(V=0,\boldsymbol{X}=\boldsymbol{x},r=\textup{Disease-unrelated reasons})},$
$\widehat{VE}_{cct}(\boldsymbol{x})$ $=1-\frac{\widehat{h}(V=1,\boldsymbol{X}=\boldsymbol{x},r=\textup{Case contact tracing})}{\widehat{h}(V=0,\boldsymbol{X}=\boldsymbol{x},r=\textup{Case contact tracing})}.$
Under our modeling assumptions and regularity conditions listed in Web Appendix B, the estimators $\widehat{VE}_{s}(\boldsymbol{x},1)$, $\widehat{VE}(\boldsymbol{x})$, and $\widehat{VE}_{cct}(\boldsymbol{x})$ converge in probability to $VE_{s}(\boldsymbol{x},1)$, $VE(\boldsymbol{x})$, and $VE_{cct}(\boldsymbol{x})$ respectively (as the sample size $n$ increases), and their normal-approximated 95% confidence intervals based on estimated SE estimators will have correct coverage in large samples. We also provide variance estimators and technical details in Web Appendix B. If $VE(\boldsymbol{x},1)$ is of interest, an additional assumption is needed to obtain $VE(\boldsymbol{x},1)=VE_{s}(\boldsymbol{x},1)$ (see Table 1), in which case $\widehat{VE}_{s}(\boldsymbol{x},1)$ will also target the estimand $VE(\boldsymbol{x},1)$.
In practice, it is common that the estimands $VE_{s}(\boldsymbol{x},1)$, $VE(\boldsymbol{x},1)$, $VE(\boldsymbol{x})$, and $VE_{cct}(\boldsymbol{x})$ are all distinct. In this case, data collected under disease-unrelated reasons and case contact tracing cannot help with the estimation of $VE_{s}(\boldsymbol{x},1)$ or $VE(\boldsymbol{x},1)$. However, the additional data are not useless: they can be used to estimate $VE(\boldsymbol{x})$, and $VE_{cct}(\boldsymbol{x})$; both estimands can be of interest depending on the nature of vaccination (intervention) and the scientific question.
Combining the stratified estimators to improve precision
In special cases where several VE estimands coincide, we are able to combine the corresponding estimators to yield a more precise estimator. If there is evidence that case contact tracing does not modify VE, we have $VE_{cct}(\boldsymbol{x})=VE(\boldsymbol{x})$ and propose a combined estimator for $VE(\boldsymbol{x})$, defined as
$\widehat{VE}_{comb}(\boldsymbol{x})$ $=w_{1}\widehat{VE}(\boldsymbol{x})+w_{2}\widehat{VE}_{cct}(\boldsymbol{x}),$
where $w_{1}$ and $w_{2}$ are arbitrary positive weights such that $w_{1}+w_{2}=1$. The weights can be simply the proportion of samples in each data source, inverse variance-weighting, or other weighting methods from the meta-analysis literature [23].
If we assume healthcare-seeking is not an effect modifier, vaccination has no direct effect on symptom alleviation, and case tracing does not modify VE, then $VE_{s}(\boldsymbol{x},1)=VE(\boldsymbol{x})=VE_{cct}(\boldsymbol{x})$, and we can construct an estimator combining all three categories of reasons:
$\widehat{VE}^{*}_{comb}(\boldsymbol{x})$ $=w_{1}^{*}\widehat{VE}_{s}(\boldsymbol{x},1)+w_{2}^{*}\widehat{VE}(\boldsymbol{x})+w_{3}^{*}\widehat{VE}_{cct}(\boldsymbol{x})$
where $(w_{1}^{*},w_{2}^{*},w_{3}^{*})$ are arbitrary positive weights with $w_{1}^{*}+w_{2}^{*}+w_{3}^{*}=1$, and these weights can be chosen similarly to $(w_{1},w_{2})$. When their assumptions hold, both combined estimators target $VE(\boldsymbol{x})$, and $\widehat{VE}^{*}_{comb}(\boldsymbol{x})$ improves precision over $\widehat{VE}_{s}(\boldsymbol{x},1)$ by leveraging the additional data. The amount of precision improvement will be greater if we have more data on asymptomatic tests. In the Covid-19 context, asymptomatic tests can often account for a large proportion of data, thereby having the potential to improve precision substantially. This result extends [13], which proposed combining stratified estimators for symptoms and case contact tracing (without specifying how to choose the weights) but did not consider disease-unrelated reasons or explain the underlying assumptions for estimation validity.
In fact, whether the above VE estimands are equal is testable from data. To evaluate whether case contact tracing modifies the VE, we can test the hypothesis $H_{0}:VE_{cct}(\boldsymbol{x})=VE(\boldsymbol{x})$ for any $\boldsymbol{x}$ such that $\widehat{VE}_{cct}(\boldsymbol{x})$ and $\widehat{VE}(\boldsymbol{x})$ can be constructed. Since both $\widehat{VE}_{cct}(\boldsymbol{x})$ and $\widehat{VE}(\boldsymbol{x})$ are normally distributed (on the root-$n$ scale) and independent in large samples, $\widehat{VE}_{cct}(\boldsymbol{x})-\widehat{VE}(\boldsymbol{x})$ also follows a normal distribution in large samples, from which $H_{0}$ can be rejected if $|\widehat{VE}_{cct}(\boldsymbol{x})-\widehat{VE}(\boldsymbol{x})|$ exceeds the critical value for a given type I error. Following a similar logic, we can also test $H_{0}:VE_{s}(\boldsymbol{x},1)=VE(\boldsymbol{x})$ using $\widehat{VE}_{s}(\boldsymbol{x},1)$ and $\widehat{VE}(\boldsymbol{x})$. This test has greater power if the candidate VEs have substantial differences, the sample sizes are large, and the VE estimates have small variances.
Simulation
In Web Appendix C, we provide simulation studies to demonstrate our results numerically. In summary, the simulation studies show that (1) the traditional aggregated odds ratio estimator can be biased for TNDs with multiple reasons for testing, (2) the proposed estimators are unbiased for the target estimands (with the necessary assumptions), (3) leveraging data for asymptomatic reasons for testing can improve precision. Furthermore, supplementary simulations to explore the effect of misclassification of the infection status and missing values in the reasons for testing are also provided in Web Appendix C.
Discussion
As TNDs become increasingly popular for evaluating VE based on observational data, the demand for valid, robust, and appropriate statistical analyses for TNDs is also growing. In this paper, we focus on an emerging issue that a TND may recruit participants with varying reasons for testing, and clarify the VE estimands associated with the various reasons. Of course, the suggested approach relies on collecting accurate information on the reason for testing at the time of recruitment participation. We note that severe bias can be caused if the reasons for testing are ignored. In summary, TNDs with various reasons for testing bring not only challenges in analyses but also present opportunities in obtaining precise estimates of VE on different populations. Not discussed here is the possibility of using estimates of $VE_{s}(\boldsymbol{x})$ and $VE(\boldsymbol{x})$ to estimate of $RR_{P}$, the relative risk for progression to symptoms associated with vaccination. Our statistical analysis and proposed methods provide a handle to take advantage of this opportunity and can potentially inform future TND studies.
In designing and implementing TND studies, it is thus important to understand why people are tested. If this information is missing, an estimated VE can be biased. In this case, the magnitude and direction of this bias is likely intractable absent further information. Therefore, when tests for asymptomatic subjects are expected, we recommend recording the reasons for testing in patient enrollment or in the surveillance system. For this purpose, an example questionnaire used by the California Department of Public Health in the United States is available in the Supplementary Material of [7].
Our statistical framework relies on the assumption that a subject should have a single, accurate reason for testing. In practice, a patient may have two or more reasons for testing or, on the contrary, unclear or misclassified reasons for testing. While this may not be an issue for mandatory testing such as case contact tracing, bias may occur with self-reporting patients. This issue can be mitigated by adopting survey methods to promote accuracy and honesty [24] and can also be examined by a sensitivity analysis. We leave this topic for future research.
To derive the VE measures given various reasons for testing, we have mainly focused on a simple setting with a binary vaccination status and a binary outcome. In the Covid-19 setting, however, the VE may differ by vaccination doses, the definition of outcomes (infected, diseased, hospitalized, or dead), and the variant of infection. With multi-level $V$ and $I$, our framework can be applied to evaluate the effect of each specific vaccination on each type of outcomes separately, provided that vaccination and outcomes are well recorded and the assumptions in Table 1 hold. On this topic, [25] proposes measuring symptom severity to alleviate selection bias on the odds ratio scale; non-binary healthcare-seeking behavior is handled as well. A comprehensive examination of these different scenarios is an important future direction.